# Tracing Nitrogen Enrichment across Cosmic Time with JWST

### v2

@@ Appendix A Optimization of strong-line diagnostic combinations for determining metallicity
To evaluate the optimal choice of diagnostics, we constructed a pool of possible line ratios and explored all combinations including at least three diagnostics. For each combination, we computed the distribution of $\chi^{2}$ values across the galaxy sample and quantified the typical goodness of fit by taking the median $\chi^{2}$. In addition, we calculated the curvature of the $\chi^{2}$ profiles as an indicator of the robustness of the metallicity determination, and we also evaluated the variance of the inferred metallicity distribution, quantified as the interquartile range $\left[P_{75}-P_{25}\right]$, to avoid solutions where all diagnostics would predict overly clustered values (thus limiting degeneracies in the inversion). During the procedure, we further excluded from the statistics those galaxies whose inferred metallicities yielded a reduced $\chi^{2}$ above a fixed threshold (e.g. $\chi^{2}_{\nu}>10$). While this inevitably decreases the final number of galaxies with indirect metallicity estimates, it ensures that the retained measurements are more robust and reliable.
@@ Appendix B The origin of the scatter in the N/O calibrations
We investigate the physical origin of the intrinsic scatter affecting the empirical N2O2–N/O and N2S2–N/O relations. In particular, we analyse the residuals from the best-fit trends as a function of key physical parameters- redshift, metallicity, and electron temperature $T_{3}$- as well as the local gradients of these quantities across the diagnostic planes. The visual trends are summarized in Fig. 8, where the three panels on the left refer to the N2O2 diagnostic and those on the right to N2S2; from top to bottom, the colour coding traces redshift, metallicity, and electron temperature.
@@
No clear or systematic trend is found with redshift. The corresponding gradient directions are nearly vertical ($\theta_{z}\approx 120^{\circ}$), indicating that any residual correlation acts primarily along the $\log(\mathrm{N/O})$ axis rather than with the diagnostic itself. To assess whether this apparent pattern could arise by chance, we performed a Monte Carlo permutation analysis in which galaxy redshifts were randomly shuffled while keeping the N/O ratios and line ratios fixed. The resulting $p$-value ($p_{\mathrm{perm}}\approx 0.07$) confirms that any dependence on redshift is statistically marginal. Consistently, the correlations between residuals and redshift are weak and not significant (N2O2: Pearson $r=0.29$, $p=0.06$; N2S2: $r=0.07$, $p=0.63$). These results indicate that the scatter in both diagnostics does not depend significantly on redshift within our sample ($z\simeq 2\text{--}3$), suggesting that the N2O2–N/O and N2S2–N/O relations are not subject to strong evolutionary effects with cosmic time.
@@
Among the parameters considered, the gas-phase metallicity shows the clearest dependence. In both diagnostics, metallicity gradients are oriented along directions intermediate between the diagnostic and the $\log(\mathrm{N/O})$ axis ($\theta_{Z}\approx-46^{\circ}$ for N2O2 and $32^{\circ}$ for N2S2), and the residuals from the best-fit relations show a strong anti-correlation with metallicity (Pearson r=0.78r=-0.78 and 0.26-0.26 for N2O2 and N2S2, respectively). Specifically, the fit tends to underestimate $\log(\mathrm{N/O})$ at low metallicity.
@@
The gradients with electron temperature are shallower and show only weak or marginal correlations with the residuals, with higher $T_{3}$ values corresponding to slightly lower N/O at fixed diagnostic ratio, going in the same direction as the trend with metallicity.
@@
The diagnostics dependence on metallicity can be interpreted as a secondary effect of the ionisation parameter $U$, which is known to anti-correlate with metallicity (e.g. Dopita & Evans, 1986; Pérez-Montero, 2014; Thomas et al., 2019). Since the [O ii] and [S ii] lines are sensitive to the ionisation structure, variations in $U$ shift the position of galaxies in both diagnostic planes: at fixed N/O, a decrease in $U$ moves galaxies toward higher N2O2 or N2S2 values. Furthermore, recent photoionisation models by Martinez et al. (2025) predict a metallicity dependence of these diagnostics, consistent with the trends observed in our data.
@@
Table 2: Summary of residual correlations and gradient orientations for the N/O calibrations. Pearson correlation coefficients refer to the residuals of the best-fit relations.
@@
N2O2 –N/O $z$ $120.4\pm 25.5$ +0.29+0.29 $0.05$
O/H $46.1\pm 0.8$ 0.78-0.78 $10^{-4}$
$T_{3}$ $140\pm 20$ +0.21+0.21 $10^{-4}$
N2S2 –N/O $z$ $70\pm 39$ +0.20+0.20 $0.11$
O/H $32.1\pm 2.0$ 0.26-0.26 $10^{-4}$
$T_{3}$ $147\pm 11$ +0.20+0.20 $10^{-4}$
Figure 8: Same as Fig. 4, but here we include six panels arranged in three rows by two columns, each showing the relation between directly measured N/O and the observed emission-line ratios [N ii]/[O $i$] (left column) and [N ii]/[S $i$] (right column). Each row adopts a different colour coding: redshift (top), gas-phase metallicity derived via the direct method (middle), and [O iii] electron temperature, $T_{3}$ (bottom).
@@ Appendix C Statistical tests on the N/O-O/H residual distributions
Table 3: Statistical comparison between local and high-$z$ N/O populations, computed both on the whole sample and in two metallicity bins 12+log(O/H) <8.1<8.1, and 12+log(O/H) $\geq 8.1$.
@@
Total sample 88,462 546 Combined: 89,008
$12+\log(\mathrm{O/H})<8.1$ 340 77
$12+\log(\mathrm{O/H})\geq 8.1$ 88,122 469
@@
Median (all) 0.924-0.924 1.063-1.063 $\Delta=-0.139$ dex
Std. deviation $0.196$ $0.261$ High-$z$ more scattered
@@
Median residual +0.01+0.01 +0.14+0.14 $\Delta=0.13$ dex
@@ Appendix D Observational biases potentially affecting the N/O abundances
To test whether detectability nonetheless influences the observed N/O–O/H distribution, we examined the S/N of the [N ii] line within our sample. However, we do not observe any clear correlation between S/N and the measured N/O values. Galaxies with strongly enhanced N/O include both weaker detections (S/N $\sim$3–5) and very secure ones (S/N >>10). This seems to indicate that detectability alone is unlikely to drive the entire observed population of N-enhanced systems. In particular, several JADES galaxies exhibit [N ii] detections with S/N >>15; even if their intrinsic [N ii] fluxes were a factor of $\sim$5 lower (i.e., consistent with a “normal” N/O ratio), they would still remain detectable within the same observations. The scarcity of such objects therefore suggests that the excess of high N/O systems is not solely an observational artifact but reflects an underlying physical trend.
@@
As visible in Fig. 9, the ratio becomes increasingly biased toward higher values as the gas density rises, and a mild vertical dependence on the ionisation parameter is also present, although it remains subdominant relative to the density effect. While it is true that real galaxies host a far more complex, multiphase ISM, the models demonstrate that very dense, unresolved substructures could have a non-negligible impact on the observed ratio.
However, robust diagnostics of such extreme densities, such as UV doublet ratios of highly ionised species (e.g., C iii], N iii], Si iii], N iv]), are not available for large high-z samples. Even when these lines are detected, it is important to note that such ions do not necessarily trace gas that is co-spatial with the regions emitting [N ii] and [O ii] (e.g. Kewley et al., 2019; Choustikov et al., 2025). In a simplified two-component picture, if the high-density component is also the brightest–i.e. it dominates the emergent line flux–then the density inferred from the UV transitions may reasonably approximate the physical conditions relevant for [N ii] as well. Conversely, if the low-density component contributes substantially to the total flux, a more sophisticated multi-phase treatment would be required to correctly assess the bias affecting N2O2. Therefore, with current data, we cannot yet precisely quantify the extent to which density effects contribute to the observed N/O enhancement.
Finally, we explored how these N-enhanced galaxies behave with respect to the star-forming main sequence, in order to assess if we are only considering starburst objects in our sample or if we are probing a fair fraction of the “normal” galaxy population. The results are shown in Fig. 11, where we also report, for reference, the MS from Popesso et al. (2023).
An inspection of Fig. 11 shows that a significant fraction of our galaxies populate the upper envelope of the main sequence, indicative of bursty or elevated star-formation activity. About $\sim 67\%$ of the systems classified as N/O-enhanced, i.e. those with $\Delta\log(\mathrm{N/O})>0.4$ relative to the local best-fit relation (see the colour coding in the figure), fall within this starburst regime. However, this trend appears quite scattered: N/O-enhanced galaxies are also found at more moderate star-formation rates, and the excess of sources above the MS at low stellar masses could still be driven by selection effects. We also performed a correlation analysis between $\Delta\log(\mathrm{N/O})$ and $\Delta MS$, obtaining $R^{2}\sim$0.1 (where $R^{2}$ represents the coefficient of determination, i.e., the fraction of variance in N/O enhancement explained by the offset from the main sequence). This indicates that $\Delta MS$ accounts for only $\sim$10% of the observed variance in N/O ratios, suggesting it is neither the sole nor the dominant driver of the N/O enhancement in our sample.
@@
Figure 11: Star formation main sequence for the high-z galaxies of our sample; the MARTA objects are marked with exagons, the JADES sample with circles and the JWST sample with diadonds. The colour coding is in $\Delta\log(\mathrm{N/O})$, i.e. the residual of log(N/O) computed from the local bestfit.

### v3

@@ Acknowledgements.
This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 1879.
Some of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST). The specific observations analyzed can be accessed via https://doi.org/10.17909/v7wt-1r79. STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555.
@@ Appendix A Optimization of strong-line diagnostic combinations for determining metallicity
To evaluate the optimal choice of diagnostics, we constructed a pool of possible line ratios and explored all combinations including at least two diagnostics. For each combination, we computed the distribution of $\chi^{2}$ values across the galaxy sample and quantified the typical goodness of fit by taking the median $\chi^{2}$. In addition, we calculated the curvature of the $\chi^{2}$ profiles as an indicator of the robustness of the metallicity determination, and we also evaluated the variance of the inferred metallicity distribution, quantified as the interquartile range $\left[P_{75}-P_{25}\right]$, to avoid solutions where all diagnostics would predict overly clustered values (thus limiting degeneracies in the inversion). During the procedure, we further excluded from the statistics those galaxies whose inferred metallicities yielded a reduced $\chi^{2}$ above a fixed threshold (e.g. $\chi^{2}_{\nu}>10$). While this inevitably decreases the final number of galaxies with indirect metallicity estimates, it ensures that the retained measurements are more robust and reliable.
We applied the same diagnostic-selection test to the local reference sample. For the low-redshift galaxies, the combination that formally yields the lowest median $\chi^{2}$ with respect to the Curti et al. (2017) calibrations is $R3$ together with $\tilde{R}$. However, the difference relative to the $R2,R3$ combination is very small, amounting to only $\sim 0.5$ in the median $\chi^{2}$ of the sample. At the same time, we find that the $R2,R3$ pair maximizes both the metallicity variance and the curvature of the $\chi^{2}$ surface, providing slightly stronger leverage in constraining the solution. Given the marginal difference in $\chi^{2}$ and in order to maintain full methodological consistency with the high-redshift analysis, we therefore adopt the $R2,R3$ combination also for the local sample.
@@ Appendix B The origin of the scatter in the N/O calibrations
We investigate the physical origin of the intrinsic scatter affecting the empirical N2O2–N/O and N2S2–N/O relations. In particular, we analyse the residuals from the best-fit trends as a function of key physical parameters- redshift, metallicity, and electron temperature $T_{3}$- as well as the local gradients of these quantities across the diagnostic planes. The visual trends are summarized in Figure 8, where the three panels on the left refer to the N2O2 diagnostic and those on the right to N2S2; from top to bottom, the colour coding traces redshift, metallicity, and electron temperature.
@@
No clear or systematic trend is found with redshift. The two gradient angles, $\sim 67^{\circ}$ for N2O2 and $\sim-14^{\circ}$ for N2S2, have large uncertainties and are therefore not particularly informative. We performed a Monte Carlo permutation analysis in which galaxy redshifts were randomly shuffled while keeping the N/O ratios and line ratios fixed. The resulting $p$-value ($p_{\mathrm{perm}}\approx 0.07$) confirms that any dependence on redshift is statistically marginal. Consistently, the correlations between residuals and redshift are weak and not significant (N2O2: Pearson $r=0.11$, $p=0.28$; N2S2: r=0.18r=-0.18, $p=0.10$). These results indicate that the scatter in both diagnostics does not depend significantly on redshift within our sample ($z\simeq 1.5\text{--}6$), suggesting that the N2O2–N/O and N2S2–N/O relations are not subject to strong evolutionary effects with cosmic time.
@@
Among the parameters considered, the gas-phase metallicity shows the clearest dependence. In both diagnostics, metallicity gradients are oriented along directions intermediate between the diagnostic and the $\log(\mathrm{N/O})$ axis ($\theta_{Z}\approx-48^{\circ}$ for N2O2 and $37^{\circ}$ for N2S2), and the residuals from the best-fit relations show a strong anti-correlation with metallicity (Pearson r=0.75r=-0.75 and 0.39-0.39 for N2O2 and N2S2, respectively). Specifically, the fit tends to underestimate $\log(\mathrm{N/O})$ at low metallicity.
@@
The gradients with electron temperature are similar but show weaker correlations with the residuals (Pearson values of +0.20 and +0.23 respectively), with higher $T_{3}$ values corresponding to lower N/O at fixed diagnostic ratio, going in the opposite direction as the trend with metallicity.
@@
The diagnostics dependence on metallicity can be interpreted as a secondary effect of the ionisation parameter $U$, which is known to anti-correlate with metallicity (e.g. Dopita & Evans 1986; Pérez-Montero 2014; Thomas et al. 2019). Since the [O ii] and [S ii] lines are sensitive to the ionisation structure, variations in $U$ shift the position of galaxies in both diagnostic planes: at fixed N/O, a decrease in $U$ moves galaxies toward higher N2O2 or N2S2 values. Furthermore, recent photoionisation models by Martinez et al. (2025) predict a metallicity dependence of these diagnostics, consistent with the trends observed in our data.
@@
Table 2: Summary of statistical tests for the N/O calibrations. [^6] Pearson correlation coefficients refer to the residuals of the best-fit relations.
@@
N2O2 –N/O $z$ $67\pm 49$ +0.11+0.11 $0.28$
O/H $48\pm 1$ 0.75-0.75 $10^{-4}$
$T_{3}$ $136\pm 14$ +0.20+0.20 $10^{-4}$
N2S2 –N/O $z$ $14\pm 26$ 0.18-0.18 $0.10$
O/H $37\pm 2$ 0.39-0.39 $10^{-4}$
$T_{3}$ $148\pm 4$ +0.23+0.23 $10^{-4}$
Figure 8: Same as Figure 4, but here we include six panels arranged in three rows by two columns, each showing the relation between directly measured N/O and the observed emission-line ratios [N ii]/[O $i$] (left column) and [N ii]/[S $i$] (right column). Each row adopts a different colour coding: redshift (top), gas-phase metallicity derived via the direct method (middle), and [O iii] electron temperature, $T_{3}$ (bottom).
@@ Appendix C Statistical tests on the N/O-O/H residual distributions
Table 3: Statistical comparison between local and high-$z$ N/O populations. [^8] Tests are computed both on the whole sample and in two metallicity bins 12+log(O/H) <8.1<8.1, and 12+log(O/H) $\geq 8.1$
@@
Total sample 88,201 664 Combined: 88,865
$12+\log(\mathrm{O/H})<8.1$ 234 121
$12+\log(\mathrm{O/H})\geq 8.1$ 87967 543
@@
Median (all) 0.923-0.923 1.067-1.067 $\Delta=+0.144$ dex
Std. deviation $0.195$ $0.226$ High-$z$ more scattered
@@
Median residual +0.001+0.001 +0.134+0.134 $\Delta=0.13$ dex
@@ Appendix D Observational biases potentially affecting the N/O abundances
To test whether detectability nonetheless influences the observed N/O–O/H distribution, we examined the S/N of the [N ii] line within our sample. However, we do not observe any clear correlation between S/N and the measured N/O values. Galaxies with strongly enhanced N/O include both weaker detections (S/N $\sim$3–5) and very secure ones (S/N >>10). This seems to indicate that detectability alone is unlikely to drive the entire observed population of N-enhanced systems. In particular, several JADES galaxies exhibit [N ii] detections with S/N >>15; even if their intrinsic [N ii] fluxes were a factor of $\sim$5 lower (i.e., consistent with a “normal” N/O ratio), they would still remain detectable within the same observations. The scarcity of such objects therefore suggests that the excess of high N/O systems is not solely an observational artifact but reflects an underlying physical trend. To further assess whether the observed N/O–O/H trend could be driven by our S/N selection on the [N ii] line, we extended the analysis to include galaxies without a formal [N ii] detection (i.e. S/N<3<3). The concern is that requiring a detection at low metallicity might preferentially select systems with intrinsically enhanced nitrogen emission, thereby artificially biasing the relation toward high N/O values.
If this were the dominant effect, we would expect that most galaxies lacking a reliable [N ii] detection would lie along the local N/O–O/H relation, with only the most extreme N-enhanced systems (i.e. those significantly above the local relation) remaining detectable. To test this scenario, we recomputed N/O and O/H for the non-detected sources using the same methodology adopted throughout the paper: direct-method abundances for the subset with robust auroral-line measurements (X objects), and strong-line calibrations for the remaining sources (Y objects). For galaxies with S/N<3<3 in [N ii], we treated the measured fluxes as $3\sigma$ upper limits and propagated them consistently into the abundance determination.
The results are shown in Figure 11. We do not observe a clear tendency for the upper limits to align with the local N/O–O/H relation. Instead, both direct- and strong-line-based upper limits populate a region broadly consistent with that defined by the detected sources, and their distribution appears compatible with the main observed trend. While this does not directly contradict our conclusions, we acknowledge that the presence of upper limits could still be hiding a lower underlying trend. For this reason, we are currently developing a forward model aimed at reproducing the “median” high-redshift galaxy population, which will be presented in forthcoming work.
As an additional check, we quantified how many of the N-enhanced galaxies would still satisfy our S/N>3>3 requirement on [N ii] if they were instead placed on the local N/O–O/H relation. For each object, we computed a “reduction factor”, defined as the factor by which its observed [N ii] flux would need to be decreased in order to match the local fit. We then evaluated whether the rescaled flux would remain above our detection threshold. We find that $\sim$43% of the galaxies would still be detected after applying this reduction. This result indicates that almost half of the enhanced sample would remain observable even if their intrinsic N/O ratios were consistent with the local relation. While this does not constitute a confirmation of our results, it at least rules out the possibility that the observed N-enhanced population is entirely driven by objects lying just at the detection limit. Furthermore, based on this result, if the local relation were the typical condition for high-z objects, we would expect to observe a non-negligible fraction of galaxies distributed along the local relation, since they would still be observable, but this is not the case.
@@
As visible in Figure 10, the ratio becomes increasingly biased toward higher values as the gas density rises, and a mild vertical dependence on the ionisation parameter is also present, although it remains subdominant relative to the density effect. While it is true that real galaxies host a far more complex, multiphase ISM, the models demonstrate that very dense, unresolved substructures could have a non-negligible impact on the observed ratio.
However, robust diagnostics of such extreme densities, such as UV doublet ratios of highly ionised species (e.g., C iii], N iii], Si iii], N iv]), are not available for large high-z samples. Even when these lines are detected, it is important to note that such ions do not necessarily trace gas that is co-spatial with the regions emitting [N ii] and [O ii] (e.g. Kewley et al. 2019; Choustikov et al. 2026). In a simplified two-component picture, if the high-density component is also the brightest–i.e. it dominates the emergent line flux–then the density inferred from the UV transitions may reasonably approximate the physical conditions relevant for [N ii] as well. Conversely, if the low-density component contributes substantially to the total flux, a more sophisticated multi-phase treatment would be required to correctly assess the bias affecting N2O2. Therefore, with current data, we cannot yet precisely quantify the extent to which density effects contribute to the observed N/O enhancement.
Finally, we explored how these N-enhanced galaxies behave with respect to the star-forming main sequence, in order to assess if we are only considering starburst objects in our sample or if we are probing a fair fraction of the “normal” galaxy population.
The results are shown in Figure 12, where we plot the offset $\Delta\log(\mathrm{N/O})$ from the local best-fit relation as a function of $\Delta_{\mathrm{MS}}$, i.e. the deviation from the star-forming main sequence of Popesso et al. (2023) computed at the redshift of each object. The black lines mark zero offsets in both quantities, while the green dashed line indicates the adopted threshold for nitrogen enhancement, $\Delta\log(\mathrm{N/O})>0.3$.
An inspection of Figure 12 shows no clear trend between $\Delta\log(\mathrm{N/O})$ and $\Delta_{\mathrm{MS}}$. Nitrogen-enhanced galaxies are distributed across the full range of offsets from the main sequence, including systems lying above, on, and below the MS. $\sim 29\%$ of the galaxies in our sample are classified as N/O-enhanced. Among these, the $\sim 32\%$ lie above the MS ($\Delta_{\mathrm{MS}}>0.3$), $\sim 39\%$ are consistent with the MS ($0.3<\Delta_{\mathrm{MS}}<0.3$), and $\sim 29\%$ fall below it ($\Delta_{\mathrm{MS}}<-0.3$). This relatively even distribution indicates that nitrogen enhancement is not confined to starburst systems, but is also common among galaxies with more typical star-formation rates.
@@
Figure 11: Same as Figure 4, but including upper limits on [N ii], shown for direct measurements (purple) and strong-line estimates (orange).
Figure 12: Representation of the offset $\Delta\log(\mathrm{N/O})$ of the high-redshift data with respect to the local $\mathrm{N/O}$ best-fit relation, as a function of $\Delta_{\mathrm{MS}}$, i.e. the offset of the same galaxies from the MS of Popesso et al. (2023), computed at the specific redshift of each object. The black dashed lines indicate the zero offsets, corresponding to perfect agreement between the observed $\log(\mathrm{N/O})$ and the local best-fit relation, and between the observed SFR and the SFR predicted by the Popesso et al. Main Sequence. The green dashed line marks the adopted enhancement threshold, corresponding to galaxies with $\log(\mathrm{N/O})$ higher by more than 0.3 dex than expected. As visible from the figure, no clear trend of the $\mathrm{N/O}$ enhancement with $\Delta_{\mathrm{MS}}$ is observed.